An insulated cup contains 75.0g of water at 24.00

o

C. A 26.00g sample of metal at 82.25

o

C

is added. The final temperature of the water

and metal is 28.34

o

C. What is the specific heat

of the metal?



1
Expert's answer
2021-05-03T07:09:20-0400

Write the equation for thermal equilibrium:


"Q_1+Q_2=0,\\\space\\
c_mm_m\Delta t_m+c_wm_w\Delta t_w=0,\\\space\\
c_m=\frac{-c_wm_w\Delta t_w}{m_m\Delta t_m},\\\space\\
\frac{-4200·75(24-28.34)}{26(82.25-28.34)}=975\text{ J/(kg°C)}."